### v1

@@ 3.2 Longitudinal Clusters of Emotions
We applied the VISTA temporal clustering approach described in the methods to the entire sample, using $M=2$ clusters for parsimony. Parameters were initialized with VISTA’s identity initialization scheme described in [31] with a latent dimension of seven, matching the seven observed features. In figure 2, we plot the noiseless trajectories of the different emotion features predicted from the estimated VISTA parameters associated to each cluster. Visually examining the noiseless trajectories, we see that the negative emotions of anger and disgust are typically higher within the second cluster (n = 3813, 31.66%), while joy is higher in the first (n = 8230, 68.34%).
@@ 4 Discussion
Predicting treatment response in real-world mental health care remains a fundamental challenge, particularly as symptom assessments are sparse or inconsistently collected. Here, we demonstrate that passively captured emotional expressions in psychotherapy transcripts, analyzed at the utterance level using a context-sensitive small language model, can identify patients at differential risk for treatment non-response. In a sample of over 12,000 psychotherapy patients we identified subgroups based on the emotions they expressed, associated these clusters with clinical outcomes, and examined emotions dynamics leveraging the psychological network approach.
@@
Results indicated that patients in the improving symptoms group exhibited flexible expression across the emotional spectrum, including joy, sadness, and neutral affect. In contrast, the non-response group displayed emotion dynamics dominated by sadness and fear, with these negative emotions exerting disproportionate influence on subsequent affective states, a pattern consistent with emotional inflexibility. The proportion of patients classified as non-responders using these passive linguistic markers aligned with estimates derived from self-report measures in prior work [2, 3], suggesting that non response can be characterized without requiring patient adherence to symptom surveys. These findings illustrate the potential of using expressed emotions as a digital marker at a large scale, especially in settings where standardized symptom assessments are inconsistently administered or incomplete. Critically, this approach requires no additional patient burden and can be implemented on existing clinical data infrastructure. By focusing on variations in how individuals express both positive and negative emotions in treatment, our findings suggest it is possible to identify affective patterns that not only predict response but also suggest personalized paths for therapeutic focus and engagement.
Our approach combines three important steps in a new way. First, we capture the dynamic trends of emotions over the course of treatment in message-based psychotherapy, similar to [8, 11, 4]. Next, we demonstrate how these trends are associated with clinical outcomes [6, 7, 9, 10]. Third, we embrace the complexity of the emotional dynamics by representing them as a network. Previous clinical studies [48, 49] suggested an association between emotion dynamics and clinical outcomes, but lacked the methodology to represent these dynamics as a network. Our method overcomes this hurdle and produces results consistent with research literature [50] demonstrating an association between emotional flexibility and a clinically significant change in both anxiety and depression symptoms.
Key to our approach was the unique contribution of VISTA [31], which uses stochastic modeling to cluster and interpret irregularly sampled psychological time series data. VISTA’s demonstrated success in clustering psychological data as measured by withheld labels over existing algorithms paved the way for its utilization in this particular setting without ground truth class membership. Our methodology is transparent at each step. VISTA parametrizes the clusters using structural equations, which enables us to use these parameters to generate networks and understand clusters’ characteristics in greater depth. VISTA easily handles data with varying numbers of observations, an important asset in analyzing psychological interventions, which can vary in duration and frequency. We make VISTA and our analytic scripts available under an open-source license, to facilitate replication and extension of these findings.
@@
Several limitations warrant consideration. First, our findings derive from message-based psychotherapy, which, although demonstrated to be equivalent in efficacy to face-to-face treatment [2], represents a specific therapeutic modality and it is not possible to determine whether our findings will generalize to other treatment settings. Future work should examine replication of our findings to non-message-based psychotherapy. Future work could also address how our analysis was split into two distinct steps (clustering and network analysis) and attempt to combine these approaches. though this would detract from the applicability of VISTA by focusing the algorithm solely on interactions among covariates rather than the current structure which allows for the generation of time series from the distributions given by the estimated parameters. Furthermore, specific linguistic results in our sample may not be representative of larger populations or in languages other than English. It will be imperative, therefore, for future research using large language models to translate from other languages to increase the sample size.
Another limitation is that our primary marker of therapy processes, emotional expression, is multiply determined [51]. Emotional expressions may represent the experience of bodily or subcortical states, for example, or appraisal of the ongoing context, schematic representation of broader situational factors, activated memories, attempts to influence the behavior of another person, or any combination of these factors [52]. Emotional communications may also be implicit or explicit [53] and subject to varying degrees of coding error [54]. Although emotional expressions in text are nonetheless heavily meaning laden and carry clear predictive value, future research should seek to move beyond emotional expressions to include other informative aspects of therapy process language.
5 Conclusion
Despite these limitations, this study demonstrates that fine-grained emotional dynamics in psychotherapy can serve as scalable, interpretable markers of treatment response. By combining AI-based SLM emotion detection with novel longitudinal clustering and network analysis, we identified distinct affective trajectories prospectively associated with clinical outcomes in over 12,000 patients. The finding that emotional flexibility, characterized by balanced expression across positive and negative emotions, distinguished responders from non-responders aligns with regulatory flexibility theory and suggests actionable therapeutic targets. Critically, our approach operates on passively collected linguistic data, circumventing patient non-adherence to survey outcomes that typically impede outcome monitoring in real-world care settings. The transparent, open-source methodology we provide offers a framework for risk stratification that could inform adaptive treatment allocation—identifying patients who may benefit from intensified intervention before clinical deterioration becomes apparent. We anticipate advances in modeling extending the usability of our methodology to different data types which can ultimately contribute to improved patient experience in personalized medicine. Future work should validate these findings across diverse therapeutic modalities, languages, and clinical populations to establish the generalizability of emotion-based digital phenotyping for precision mental health care.
@@ 7 Supplementary Material
In supplemental table 1, we report the Bonferroni-corrected p-value [42] for each individual symptom, using the Mann-Whitney U test [41] to determine if the distribution of symptoms in each of the clusters returned by VISTA differs, using both the first and last symptoms available for each patient during the study period.

### v2

@@ 3.2 Longitudinal Clusters of Emotions
We applied the VISTA temporal clustering approach described in the methods to the entire sample, using $M=2$ clusters for parsimony. Parameters were initialized with VISTA’s identity initialization scheme described in [brindle2024vista] with a latent dimension of seven, matching the seven observed features. In figure 2, we plot the noiseless trajectories of the different emotion features predicted from the estimated VISTA parameters associated to each cluster. Visually examining the noiseless trajectories, we see that the negative emotions of anger and disgust are typically higher within the second cluster (n = 3813, 31.66%), while joy is higher in the first (n = 8230, 68.34%).
@@ 4 Discussion
Predicting treatment response in real-world mental health care remains a fundamental challenge, particularly when symptom assessments are sparse or inconsistently collected. Here, we demonstrate that passively captured emotional expressions in patient utterances, analyzed using a context-sensitive language model, can identify differential risk for treatment non-response. In a sample of over 12,000 psychotherapy patients we identified subgroups based on their expressed emotions, associated these clusters with clinical outcomes, and examined emotions dynamics using network analysis.
@@
Patients with improving symptoms exhibited flexible expression across the emotional spectrum, including joy, sadness, and neutral affect. In contrast, the non-response group displayed dynamics dominated by sadness and fear, with these negative emotions exerting disproportionate influence on subsequent affective states. These patterns align with the regulatory flexibility framework [bonanno2013regulatory, bonanno2021resilience]: adaptive regulation involves balanced, context-sensitive expression, while inflexibility manifests as rigid, negatively-biased patterns. Notably, the proportion of patients classified as non-responders using these passive linguistic markers aligned with estimates derived from self-report measures in prior work [pullmann2025message, hull2020two], suggesting that non response can be characterized without requiring patient adherence to symptom surveys.
Our approach combines extends prior work [Shapira2022, Eberhardt2025, malgaroli2023natural, Burkhardt2021, Ewbank2021, Malgaroli2024, Nook2022] in three ways: capturing emotion at utterance-level granularity, linking temporal clusters to outcomes, and interpreting their emotion dynamics using network analysis. VISTA-SSM [brindle2024vista] enables this by clustering irregularly sampled time series via stochastic modeling, parameterizing clusters with structural equations that can be used to generate networks and better understand their characteristics. Previous clinical studies [coifman2007affect, heininga2019dynamical] suggested associations between emotion dynamics and clinical outcomes, but lacked the methodology to represent these dynamics as a network. Our method overcomes this hurdle and produces results consistent with research literature demonstrating associations between emotional flexibility and a clinically significant change in both anxiety and depression symptoms.[kuppens2017emotion]. We make VISTA-SSM and our analytic scripts available under an open-source license, to facilitate replication and extension to different data types.
@@
Several limitations warrant consideration. First, our findings derive from message-based psychotherapy, which, although demonstrated to be equivalent in efficacy to face-to-face treatment [pullmann2025message], represents a specific therapeutic modality. Future work should examine replication of our findings to other delivery modalities. Future work could also address how our analysis was split into two distinct steps (clustering and network analysis) and attempt to combine these approaches. though this would detract from the applicability of VISTA-SSM by focusing the algorithm solely on interactions among covariates rather than the current structure which allows for the generation of time series from the distributions given by the estimated parameters. Furthermore, specific linguistic results in our sample may not be representative of larger populations or in languages other than English. It will be imperative, therefore, for future research using language models to include other languages to increase generalizability of findings. Another limitation is that our primary marker of emotional processes, emotional expression, is multiply determined [lange2020toward]. Emotional expressions may represent the experience of bodily or subcortical states, for example, or appraisal of the ongoing context, schematic representation of broader situational factors, activated memories, attempts to influence the behavior of another person, or any combination of these factors [ledoux2017higher]. Although emotional expressions in text are nonetheless heavily meaning laden and carry clear predictive value, future research should seek to include other emotional features.
This study demonstrates that fine-grained emotional dynamics can serve as scalable, interpretable markers of treatment response, operating on passively collected data without patient burden. Critically, this approach requires no additional patient burden and can be implemented on existing clinical digital infrastructure. Emotion flexibility offers an interpretable framework for risk stratification that could identify patients needing intensified intervention before clinical deterioration. Future work should validate these findings across diverse therapeutic modalities, languages, and clinical populations to establish the generalizability of emotion-based digital phenotyping for precision mental health care.
@@ 6 Supplementary Material
In supplemental table 1, we report the Bonferroni-corrected p-value [bonferroni1936teoria] for each individual symptom, using the Mann-Whitney U test [mann1947test] to determine if the distribution of symptoms in each of the clusters returned by VISTA differs, using both the first and last symptoms available for each patient during the study period.